OpenStudy (anonymous):

Determine the intervals on which f(x)=sqrt((x+1)/x) is continuous

OpenStudy (anonymous):

\[f(x)=\sqrt{\frac{x+1}{x}}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this function, like almost all that you know, is continuous on its domain, so what you are really being asked for is the domian of this function. that boils down to finding where the expression inside the square root is greater than or equal to zero. of course you have to exclude zero as well, because you may not divide by zero

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

i.e. you need to solve \[\frac{x+1}{x}\geq 0\]

OpenStudy (anonymous):

is that ok?

OpenStudy (anonymous):

I think so, wouldn't it be 1>infinity

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\[\frac{x+1}{x}\geq 0\] this thing changes sign at both 0 and 1 it is postive if \(x<-1\) or if \(x>0\)

OpenStudy (anonymous):

sorry i meant changes sign at \(x=-1\) and at \(x=0\)

OpenStudy (anonymous):

i think i found out is it (-infinity,-1)U(0,infinity)

OpenStudy (anonymous):

yes, that is it although the -1 should be included

OpenStudy (anonymous):

well maybe not because it would not be continuous from the right, so stick with your answer

OpenStudy (anonymous):

oh okay thank you

OpenStudy (anonymous):

your answer is correct

OpenStudy (anonymous):

yw